### v1

# Faraday Depolarization Study of a Radio Galaxy Using LOFAR Two-metre Sky Survey: Data Release 2

Abstract
We present a detailed depolarization analysis of a radio galaxy ILTJ012215.21+254334.8, utilizing polarimetric data from the LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) catalogue. The selected source exhibits a rotation measure (RM) of rad m-2 and a projected linear size of kpc at a redshift . Depolarization model fitting was performed on LOFAR High Band Antenna data (120–168 MHz), with fractional polarization detected at . Five depolarization models were tested, and Bayesian qu-fitting revealed that the three-component model (1T+2ED) best describes the data, with a reduced chi-squared value of 2.12 and a logarithmic Bayesian evidence of 1384.82. This model includes a Faraday-thin component at RM rad m-2 (instrumental leakage) and two external Faraday dispersion astrophysical emission at RM rad m-2. The results demonstrate that depolarization in low-frequency radio galaxies requires multi-component modelling and is driven by turbulence and inhomogeneity in the magneto-ionic medium. Our findings highlight the potential of LOFAR polarization studies for probing galactic and intergalactic magnetic fields with high precision.
keywords:
techniques: polarimetric, galaxies: active, galaxies: magnetic fields, radio continuum: galaxies, radiation mechanisms: non-thermal1 Introduction
Radio galaxies are energetic extragalactic objects distinguished by vast radio-emitting lobes that extend well beyond the optical extent of their host galaxies. These lobes are powered by highly collimated plasma jets launched from an active galactic nucleus (AGN), itself driven by accretion onto a central supermassive black hole (SMBH) (Oei et al., 2024). Their radio luminosities can reach up to across frequencies from 10 MHz to 100 GHz, primarily due to synchrotron radiation emitted by relativistic electrons spiralling around magnetic field lines. Typically hosted by massive elliptical galaxies, the term radio galaxy refers both to the host galaxy and its extended jet-lobe system. The SMBH at the core powers the jets through accretion of surrounding material (Hardcastle and Croston, 2020). The lobes, often observed as large, symmetrical pairs, represent the most extended structures, sometimes spanning sizes of order megaparsecs (Oei et al., 2023). Their radio emission is frequently highly linearly polarized (Begelman et al., 1984; Miley, 1980).
Radio galaxies exhibit a wide range of physical sizes, from a few parsecs up to several megaparsecs, and are morphologically classified according to the Fanaroff-Riley (FR) scheme (Fanaroff and Riley, 1974). FR II galaxies are powerful, edge-brightened sources with luminous hotspots at the extremities of their lobes. They commonly show weak but well-collimated jets that often appear one-sided owing to relativistic beaming, with radio luminosities typically above at 178 MHz (Fanaroff and Riley, 1974). These systems are typically associated with quasars or more distant radio-loud galaxies. Conversely, FR I galaxies are weaker, edge-darkened sources whose emission is brightest near the core and decreases outward. They generally lack prominent hotspots, and their jets and lobes often appear distorted. With radio luminosities below , FR I sources are more frequently found in nearby galaxies or dense galaxy clusters environment. Although, Mingo et al. (2019) recently confirmed for LoTSS catalogue radio luminosity is not always a reliable way to determine if a source is edge-brightened (FR-II) or centre-brightened (FR-I).
Extragalactic foreground radio sources have played a crucial role in advancing our understanding of the cosmic evolution of magnetic fields in the Milky Way, in external galaxies, and in the large-scale cosmic web (e.g., Taylor et al. (2009), Harvey-Smith et al. (2011), Mao et al. (2010)). The study of radio polarization provides further insight into the morphology of radio galaxies and their surrounding environments (O’Sullivan et al., 2012; Pasetto et al., 2016). Broadband radio spectro-polarimetric observations (Farnsworth et al., 2011; Anderson et al., 2015; O’Sullivan et al., 2017; Pasetto et al., 2018) have proven particularly effective for investigating the magneto-ionic medium in AGNs by measuring Faraday rotation measures (RMs).
When synchrotron radiation propagates through a magneto-ionic medium, or Faraday screen, the polarization angle experiences a rotation known as Faraday rotation (FR). Simultaneously, the polarized flux density may be reduced through processes collectively referred to as depolarization. Depolarization is generally divided into two main categories: external depolarization and internal depolarization (Pasetto et al., 2018). Modelling these effects requires considering the structure of the magnetic field, which may range from fully ordered to a mixture of ordered and turbulent components (Burn, 1966; Sokoloff et al., 1998; Rossetti et al., 2008; Yushkov et al., 2024).
In this paper, we investigate the depolarization properties of a single source drawn from the LoTSS DR2 catalogue (O’Sullivan et al., 2023). Located at a redshift of 0.05, with a radio luminosity of at 144 MHz, this source can be a FR I radio galaxy (Fanaroff and Riley, 1974). We employ the qu-fitting technique to model the fractional Stokes parameters as functions of wavelength squared. Recent studies have demonstrated that this approach is often more sensitive than RM synthesis for detecting multiple Faraday components (Farnsworth et al., 2011; O’Sullivan et al., 2012; Sun et al., 2015).
The study of Faraday depolarization in radio galaxies provides a critical window into the magneto-ionic environments both within and surrounding these powerful extragalactic sources (Strom, 1973). Leveraging the LOFAR Two-metre Sky Survey Data Release 2 (LoTSS-DR2), which offers high-sensitivity, high-resolution polarization data at low radio frequencies enables unprecedented insight into the interplay between magnetic fields, plasma conditions, and large-scale radio structures. The paper is organized as follows: Section 2 introduces the key observables of radio polarization and Faraday rotation; Section 3 describes the depolarization models applied in our analysis; Section 4 outlines the main properties of the selected LoTSS DR2 source; Section 5 details the methodology; Section 6 presents the results; and Section 7 discusses the implications and provides prospects for future work.
2 Polarization and Faraday Rotation
Polarization in radio galaxies offers direct evidence for synchrotron radiation, caused by relativistic electrons spiralling along magnetic fields—which naturally produces linearly polarized light. Measuring the degree and orientation of polarization reveals the strength and structure of magnetic fields in jets and lobes, as well as information about the galaxy’s orientation relative to the observer.To quantitatively describe the polarization state of electromagnetic radiation, a set of four values is utilized known as Stokes parameters: I, Q, U, and V. These parameters are experimentally convenient as they correspond to sums or differences of measurable intensities, allowing for direct determination from observations. The complex linear polarization is defined as:
here, I, Q, and U are the measured Stokes parameters, and p and are the fractional polarization and the polarization angle of the polarized wave, respectively, described as :
Here, is the polarized intensity, and as the fractional values of the Stokes parameters respectively, and .
The study of the polarization states of radio emissions enables us to analyse two important effects: Faraday Rotation and Faraday Depolarization (Anderson et al., 2015; Piras, 2024). Faraday Rotation is caused by the rotation of the plane of polarization of electromagnetic waves as they journey across vast cosmic distances interacting with magnetized medium. Crucially, the extent of this rotation directly encodes vital information about the magneto-ionic medium—the ionized gas and embedded magnetic fields. This makes Faraday rotation an exceptionally powerful tool for probing magnetic fields not just within galaxies, but also in the vast and tenuous intergalactic medium, providing unique insights into the large-scale magnetization of the Universe that cannot be obtained through other means (Brentjens and de Bruyn, 2005). To accurately characterize the Faraday rotation for these distinct regions of polarized emission, the concept of Faraday Depth is utilized. It is specifically defined as the integral of the thermal electron density (in units of )and the line-of-sight component of the magnetic field (in ) over the physical distance (x) (in parsec) to the source, typically expressed as (Pasetto et al., 2018) :
In the most straightforward scenario, where a polarized radio source lies behind a single, uniform layer of magnetized plasma, the concept of Faraday depth simplifies considerably: it becomes equivalent to the Rotation Measure (RM) itself. In such an ideal case, the observed polarization angle is directly determined by the source’s intrinsic polarization angle and the Faraday rotation (RM), following the simple linear relationship (Anderson et al., 2016):
Here, represents the square of the observing wavelength, illustrating how the rotation of the polarization plane is directly proportional to both the RM of the intervening medium and the square of the wavelength at which the observation is made (Burn, 1966; Brentjens and de Bruyn, 2005).
3 Depolarization
Depolarization occurs because variations in the magnetic field or electron density along the line of sight, or across the telescope beam, cause the polarization angle to rotate by different amounts in different regions, reducing the net observed polarization (Burn, 1966; Goodlet et al., 2004). When these diverse polarization orientations are averaged across the observing beam, they tend to cancel each other out, diminishing the overall polarization signal. Hence, two types of depolarization come into play: internal and external (Pasetto et al., 2018). The remaining polarization percentage is effectively utilized to gauge out the degree of order and orientation within the magnetic field, whether it resides in the radio source itself or in the intervening inter galactic medium (IGM). The complex polarized signal observed from radio sources is fundamentally described by a model that accounts only for Faraday rotation and does not depolarize the signal is given as (Burn, 1966) :
Here, represents the intrinsic fractional polarization of the radiation, and is the intrinsic polarization angle at the source of the emission. The term describes the Faraday rotation caused by the foreground magneto-ionic medium, where RM is the Rotation Measure and is the square of the observing wavelength. This model is also known as ‘Faraday Thin’ Component (Burn, 1966). In real scenarios, the magnetic field is not perfectly ordered, there exists random and uniform fields causing depolarization (Goodlet et al., 2004). The various depolarization contributions are thus added to this fundamental equation 5.
When the polarized radiation passes through an external magneto-ionic medium with a turbulent magnetic field, the depolarization caused is called ‘External Faraday Dispersion’. This is modelled as External Depolarization Component and is represented by the following equation :
Here, is the Faraday dispersion about the mean RM across the radio source on the sky.
In a specific scenario, where the region producing the synchrotron radiation and the area causing the Faraday rotation are co-spatial, the resulting depolarization is internal, referred to as ‘Differential Faraday Rotation’ (Sokoloff et al., 1998). However, in this particular model, the magnetic field is assumed to be purely regular, meaning any turbulent magnetic field component is neglected. Under these conditions, radiation from different depths within the source will undergo varying amounts of Faraday rotation, as the path length through the magnetized medium differs for each region. This is modelled as a ‘Faraday Thick’ Component. It is represented by the equation :
Here, the R is the Faraday Depth along the region and the observable RM is equal to in this case. Furthermore, depolarization can also arise when there is a gradient in the Faraday depth (O’Sullivan et al., 2017) across the source itself, or within a foreground screen that is local to the emitting region, even if the magnetic field within that region is considered uniform. In such instances, the Faraday depth effectively takes the form of a Rotation Measure (RM) that varies across the telescope’s observing beam. This spatial variation in RM causes different parts of the signal within the beam to have distinct polarization orientations, which then tend to cancel each other out when averaged, resulting in observed depolarization. In general, depolarization towards longer wavelengths can result from a combination of factors, including the mixing of the emitting and rotating media, as well as the limited spatial resolution of our observations (for more details see Sokoloff et al. (1998)).
4 The LoTSS DR2 Catalogue and the Selected Source
The LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) has compiled a comprehensive catalogue of 2,461 extragalactic sources with high-precision Rotation Measure (RM) values, covering 5,720 square degrees of the sky (O’Sullivan et al., 2023). This extensive dataset, with a polarized source density of approximately 0.43 sources per square degree, was constructed by deriving linear polarization and RM properties using RM synthesis from Stokes Q and U channel images. Observations were conducted at an angular resolution of 20 arcseconds, spanning a frequency range of 120 to 168 MHz with a channel bandwidth of 97.6 kHz. The study found that roughly of total intensity sources (brighter than 1 mJy beam-1) exhibited polarization, with a median detection threshold of 0.6 mJy beam-1 and a median RM uncertainty of 0.06 rad m-2 (though a systematic uncertainty of up to 0.3 rad m-2 remains possible after ionospheric correction). The detected sources showed a median degree of polarization of , ranging from to (O’Sullivan et al., 2023). The LoTSS DR2 RM Grid Catalogue primarily contains Fanaroff-Riley Class II (FR-II) objects, with fewer than 10% classified as Fanaroff-Riley Class I (FR-I) sources. With a luminosity cut-off of W Hz-1, there are approximately 222 sources in the LoTSS-DR2 catalogue, most of which are expected to be FR-Is (Fanaroff and Riley, 1974). However, recent study shows (Mingo et al., 2019), there can be a large population of low-luminosity FR-IIs below the expected FR break. Hence, the relative population of FR-I and FR-IIs may differ.
In this pilot study, we focus our depolarization modelling on a source, ILTJ012215.21+25433.8, with an estimated luminosity of W Hz-1 at 144 MHz, RM of rad m-2, and an estimated projected linear size of 334.94 kpc (O’Sullivan et al., 2023). A relatively larger projected linear size in radio galaxies like ILTJ012215.21+254334.8 enables more reliable and physically informative depolarization modelling, as the increased path length through the magnetized plasma allows more lines of sight for Faraday rotation and depolarization effects to manifest. Furthermore, the source’s projected linear size is close to the LoTSS-DR2 median linear size of kpc (O’Sullivan et al., 2023), providing an opportunity to investigate depolarization properties of a typical galaxy in the LoTSS-DR2 catalogue. In a subsequent study, we plan to investigate the depolarization properties of all FR-I samples in the DR2 catalogue.
A few key parameters of the selected source are listed in Table 1.
| Rotation Measure (RM) | rad m-2 |
| Polarized Intensity (P) | mJy beam-1 |
| Fractional Polarization () | 2.99 % |
| Total Intensity (I) | mJy beam-1 |
| Radio Luminosity at 144 MHz () | W Hz-1 |
| Linear Size () | 334.94 kpc |
| Redshift (z) | 0.0544 |
Here, RM is the Rotation Measure; P is the linear polarization; is the fractional polarization; I is the total intensity; is the luminosity at 144 MHz; is the projected linear size; and z is the redshift.
5 Methods
5.1 Depolarization models
When analysing LOFAR polarization measurements, it is essential to account for instrumental polarization, which typically appears as a component centered near a Faraday depth of in the Faraday spectrum. Due to ionospheric variations, this instrumental polarization can spread across a range of Rotation Measure (RM) values (Jelić et al., 2014; Mevius, 2018). Therefore, this effect must be carefully modelled; otherwise, it may contaminate the detection of genuine polarized emission.
To accurately fit the observed Stokes and values for the selected source, we considered both single and two-component models, namely: Faraday-thin (T), Faraday-thick (Tk), and external Faraday dispersion (ED), as well as combinations of these. The three model types are described as follows:
where, is the variable and are the parameters. These models are combined in the five models that are listed in the table 2 given below (Piras, S. et al., 2024). We use them to find out the best fit depolarization model that fits the observed q, u spectra for the source.
| Model | Description | Equations |
| 1ED | 1 External Faraday Dispersion | 9 |
| 1T+1ED | Faraday Thin + 1 External Faraday Dispersion | 8 + 9 |
| 1T+2ED | Faraday Thin + 2 External Faraday Dispersion | 8 + 9 + 9 |
| 1Tk+1ED | Faraday Thick + 1 External Faraday Dispersion | 10 + 9 |
| 1Tk+2ED | Faraday Thick + 2 External Faraday Dispersion | 10 + 9 + 9 |
5.2 The priors
In qu-fitting, setting good starting assumptions (priors) for the model parameters is important because they help guide the fitting process in the right direction. The selection of these priors is often informed by the observed Faraday spectrum of the source (Piras, 2024). For example, when combining multiple models, the sum of the initial polarisation percentages for all components must not exceed 1, and each initial polarization angle is limited between 0 and 180 degrees for all the models considered in Table 2.
Specific constraints are applied based on the nature of the depolarization models:
Instrumental polarization, which often appears near Faraday depth 0 in LOFAR polarization observations, also requires specific prior constraints. Instrumental polarization can manifest across a range of RM values because it arises from RM fluctuations caused by the ionosphere throughout the observation (Mevius, 2018). The instrumental polarization leakage is modelled as a Faraday-thin or top-hat Faraday thick (Burn-slab) component with its Rotation Measure confined near zero ( to for Faraday thin and to for thick components respectively), capturing the typical RM range where leakage occurs in LOFAR data (O’Sullivan et al., 2023). We adopt a uniform prior to constraint the instrumental polarization. This method allows for the separation of intrinsic source polarization from instrumental artifacts, enhancing the accuracy and reliability of the resulting Faraday rotation measures and polarization parameters.
For models involving a single external Faraday dispersion (1ED), the Rotation Measure (RM) is allowed to vary within the range of the LoTSS catalogue RM values (O’Sullivan et al., 2023). Here, we also used an uniform prior.
If the model includes two external Faraday dispersion components (2ED), the uniform prior range is chosen such that one RM component is constrained within of the LoTSS catalogue RM, while the other is allowed a broader range, for example, from to . Additionally, the width of the Rotation Measure, , is forced to be positive.
5.3 qu-fitting
To analyse the observed Stokes and values, along with their errors ( and ), a one-dimensional qu-fitting technique was applied to the LoTSS DR2 polarization spectra of the source111https://lofar-mksp.org/data/. The Stokes spectral behaviour for a 97.6 kHz channel bandwidth is unreliable and thus not included in the spectra (O’Sullivan et al., 2023). However, we generate model Stokes values using the 144 MHz value, assuming a spectral index of -0.7 for the source. For the Stokes error (), we estimate the uncertainty across frequency by assuming random variations on the single-frequency catalogue value. These values are then used to create the fractional and spectra. The corresponding and errors ( and ) are calculated as,
We used the RM-TOOLS software222https://github.com/CIRADA-Tools/RM-Tools/tree/master (Purcell et al., 2020), which incorporates the MultiNest algorithm (Feroz and Hobson, 2008; Feroz et al., 2009, 2019). MultiNest, implemented via PYMULTINEST (Buchner, J. et al., 2014), is employed to efficiently explore the parameter space of the chosen model, identify the best-fit parameters, and estimate their associated uncertainties. The algorithm provides the posterior distribution for the model parameters and the Bayesian evidence () for the model, which quantifies the probability of observing the data given a specific model.
To compare and select between different models (e.g., model “a” versus model “b”), the Bayes odds ratio is used as a metric for model selection. The Bayes odds ratio is defined as:
Here, and are the Bayesian evidences of models a and b, respectively. To evaluate the performance of different models in qu-fitting, we follow established statistical guidelines, such as those proposed by Kass and Raftery (1995), which rely on the Bayes odds ratio . By computing the odds ratio between the model exhibiting the highest Bayesian evidence and all other candidate models, the model with the lowest resulting Bayesian odds ratio value is identified as the best-fitting representation of the observed data.
In addition to the Bayesian odds ratio, the qu-fitting software also calculates the reduced chi-squared for each model. This metric assesses the goodness of fit, taking into account the model’s complexity. It is defined as:
Here, DoF represents the number of degrees of freedom, is the number of data points (equivalent to the effective number of frequency channels), and are the values predicted by the model, and and are the uncertainties associated with q, u respectively (Piras, S. et al., 2024).
6 Results
The comparison of different models based on the odds ratio and the reduced chi-squared statistic () indicate that the best description of the depolarization behaviour of the source is provided by the 1T+2ED model. This particular model includes three polarization components: one Faraday-thin term corresponding to instrumental leakage, located at a rotation measure (RM) of about , and two external depolarization (2ED) components, centred near , which likely represent the astrophysical polarized emission (O’Sullivan et al., 2023). The astrophysical emission is better represented by the superposition of two external depolarization components at similar RMs, each capturing slightly different aspects of turbulence or inhomogeneities in the magneto-ionic medium along the line of sight (Sokoloff et al., 1998; O’Sullivan et al., 2017).
The fit is supported by both the Bayesian odds ratio, which quantifies the relative probability of competing models, and the , which indicates the goodness-of-fit compared to observational uncertainties (Table 3).
| Model | N | N | DoF | ln(Z) | 2ln(Z) | |
| 1T+2ED | 11 | 3 | 814 | 2.12 | 1384.82 | 0 |
| 1Tk+2ED | 11 | 3 | 814 | 2.12 | 1376.30 | 17.05 |
| 1T+1ED | 7 | 2 | 818 | 2.77 | 1154.81 | 460.04 |
| 1Tk+1ED | 7 | 2 | 818 | 2.77 | 1148.50 | 472.65 |
| 1ED | 4 | 1 | 821 | 2.94 | 1097.36 | 574.93 |
Table 4 presents the best-fitting parameters of the models, where the rows are sorted by 2ln(Z) values.
| Model | (%) | RM(1,2,3) (rad m-2) | (rad m-2) | R (rad m-2) | (deg) |
| 1T+2ED | — | — | |||
| 1Tk+2ED | — | — | |||
| 1T+1ED | — | — | |||
| 1Tk+1ED | — | — | |||
| 1ED | — |
Figure 1 shows the observed fractional Stokes parameters q, u, and the polarized fraction as functions of wavelength squared. The over plotted curves correspond to the best-fit parameters of the 1T+2ED model. Fractional Stokes parameters q, u generally exhibit sinusoidal or oscillatory variations due to Faraday rotation, but with decreasing amplitude as increases. This is a hallmark of depolarization over wider wavelength ranges, polarization is partially cancelled due to the presence of multiple Faraday rotating regions or turbulent RM dispersion, causing the coherent addition of polarization vectors to diminish. We find the polarized fraction versus here show a hint of monotonic decline as increases beyond a certain point. This is because increasing wavelength accentuates differences in RM along the line of sight or within the beam, causing more depolarization at longer wavelengths.
Figure 2 shows the polarization angle as a function of wavelength squared. This plot reveals the rotation of the plane of polarization due to the Faraday effect. We observe an overall linear increase or decrease of with , indicating a coherent Faraday rotation caused by a relatively uniform magnetized plasma along the line of sight. However, the best-fit model also indicates some flattening or irregular fluctuations around , suggesting that partial depolarization is occurring around these wavelengths.
Figure 3 presents the polarization vector in the complex plane, where the horizontal axis is q and the vertical axis is u. Ideally, without depolarization, the data would trace a circular arc of constant radius as changes. However, in the observed data the radius shrinks, showing that the polarization fraction decreases with increasing . This contraction of the polarization vector’s length is direct evidence of depolarization, and the best-fit model captures this trend adequately. Moreover, higher frequency data (red) near the centre suggest strong and coherent polarization, while lighter points (violet/blue) that spread outward indicate depolarization and more complex polarization behaviour at lower frequencies. This pattern is characteristic of Faraday depolarization in polarized radio sources. Finally, the posterior distribution of the best-fit model parameters along with error bars is shown in Figure 4. Note that the derived values of in Table 4 are not meaningful due to the lack of absolute calibration for the polarization angles.
7 Summary and Conclusion
In this study, we have investigated the depolarization properties of a radio galaxy ILTJ012215.21+254334.8 using polarization data from the LOFAR Two-metre Sky Survey (LoTSS) DR2 catalogue. Polarization analysis provides direct insight into both the magnetic field structure and the magneto-ionic environment along the line of sight, offering a diagnostic tool for probing astrophysical jets, lobes, and the intervening medium.
To model the observed polarization behaviour, we employed a suite of depolarization models - including Faraday-thin, Faraday-thick, and external Faraday dispersion components - and combined them in a Bayesian qu-fitting framework. Statistical comparison using Bayesian evidence and reduced chi-squared metrics identified the three-component 1T+2ED model as providing the best fit to the data. This model consists of a Faraday-thin component interpreted as instrumental polarization leakage (near zero RM) and two external Faraday dispersion components centred at , consistent with values derived from the LoTSS DR2 catalogue. A relatively higher reduced chi-squared () value of 2.12 suggests that the errors have been underestimated, or that there are systematic discrepancies unaccounted for by the model. Pinpointing the exact cause of the difference is difficult because it is probably due to both the model not capturing all details and the noise or errors being underestimated. The discrepancy likely arises from a combination of these two factors.
The necessity for multiple external depolarization components suggests a complex magneto-ionic medium with significant turbulence and inhomogeneities. Direct observational evidence for depolarization includes the contraction of the polarization vector radius in the - plane and a hint of monotonic decline in polarized fraction as a function of wavelength squared. Posterior analysis further supports the role of both turbulence and line-of-sight structure in shaping the observed polarization. Selecting a source with a relatively large projected linear size (335 kpc) has enabled a robust investigation of depolarization effects in a typical low-frequency radio galaxy. The polarized intensity, RM properties measured are consistent with the LoTSS DR2 sample.
In conclusion, our results demonstrate the efficacy of low-frequency polarization observations, combined with quantitative qu-fitting, in separating instrumental effects from astrophysical depolarization and Faraday rotation signatures. The best-fitting model reveals the importance of multi-component depolarization in a typical radio galaxies of the LoTSS DR2 catalogue, reflecting the inherent complexity of their magnetized environments, usually found in dense cluster environments. In future, extending this analysis to the full FR-I sample in the LoTSS DR2 catalogue, will help establish broader statistical trends and further constrain the interplay between turbulence, environmental density, source size, and polarization properties in extragalactic radio sources.
Acknowledgments
SSP and AG would like to express their gratitude to Banwarilal Bhalotia College, Asansol, under Kazi Nazrul University, for facilitating the M.Sc. Physics program. This research forms part of the M.Sc. thesis work conducted in the final semester. The authors appreciate ChatGPT for helping with AI-driven copy editing and improving the manuscript’s clarity and style. AG acknowledges the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board (SERB), a statutory body under the Department of Science and Technology (DST), Government of India. AG would also like to thank IUCAA, Pune, for their support through the Associateship Programme and for providing access to their computational facilities.
Appendix A Fit results for 1Tk+2ED model
In this appendix, we present the fit results for the next best-fit depolarization model, designated as the model. The following plots illustrate the polarization fraction (p, q, u), polarization angle as functions of wavelength squared, alongside the modelled curves derived from the best-fit parameters (Figure 5, and 6). We also include the posterior distribution of the model parameters which are displayed through one and two-dimensional projections in a corner plot (Figure 7).